two groups of tourists traveled by automobiles. One party averaged 45 kilometers per hour and the other 35 kilometers per hour. The faster party traveled 30 km farther than the slower party and made the trip in two hours less time. How long did each party travel?
Answer by mananth(16946)

let the distance they travel be x km
speed of one party = 35 kmph
speed of other party = 30 kmph
..
x/30 -x/35 = 5/2
(35x-30x) / 30*35 = 5/2
5x/1050 =5/2
5x= 1050*5/2
5x= 525*5
x= 525 distance
..
525 /30 = 17.5 hoours
525/35= =15 hours
...
m.ananth@hotmail.ca